Compton Meeting, June 9, 2011

From PolWiki
Jump to: navigation, search

Previous Meeting: Compton Meeting, June 2, 2011

Following Meeting: Compton Meeting, June 16, 2011



  1. Updated schedule
    • Laser pre-alignment change request being written: move to early August


  1. Replay status
    • submission scripts tested and integrated in batch manager (see [1])
    • need all offline macros self-contained in svn, only what is in svn will be used in replay. Initial import done but untested.
    • --Wdconinc 16:08, 7 June 2011 (UTC)

Photon Detector

  1. A tale of 2 runs - large backgrounds and their subtraction


What information from the Compton parameters do we want to go into the database ?

(The book-keeping currently done in 'run_list' files can be automated by using the database)


Attendance: Manos, Don, Juan Carlos, Mark, Vladas, Seth, Amrendra, Jeff, Wouter, Dave

  • SAD schedule
    • moving survey & alignment to later, first/second week of August, messing with laser alignment earlier, change request in with Diane Napier after mess-up in unannounced scheduling meeting while Dave on vacation
    • start messing with laser on June 27
      • measure transfer function
      • new bellows, new cavity mirrors
      • machining of gimbal mounts (might be hard, has been in beam line, check with radcon)
      • put everything back before Don goes on vacation
    • September 19: re-fiducialization
    • September 30: final alignment
    • somehow vacuum should be pulled between these two dates (disappeared from schedule)
  • Laser
    • new cavity mirrors, transfer function measurement with new mirrors
    • continued work on laser polarization measurement by Seth
  • Photon detector
    • got approval for translation stage for Compton cone position measurement
  • Analysis
    • batchman scripts work now
    • only 1TB work disk, not enough room for electron detector (no ROOT analyzer yet)
    • Dave's suggestion: copy things over to cdaq cluster while replay is running
    • full replay to tape silo, jget subset to work disk
    • changes in data acquisition should not affect replay continuity
    • some changes in ROOT tree names, will be sent out by email
  • Tale of two runs
    • ELOG Photon 121
    • asymmetry changed when adding cuts
    • fixed width histograms, large effect when opening up range
    • yield for background subtraction, change in average asymmetry
    • Don/Dave will rerun scripts on larger set of runs to get uncertainty
  • Electron detector
    • ELOG Electron 217
    • dependence of polarization on photo-cathode QE
    • QE also depends on beam current, so lowest QE point could have large horizontal uncertainty
    • link to wiki